# AirIO: Learning Inertial Odometry with Enhanced IMU Feature Observability air-io.github.io

### v1

@@ Abstract
Inertial odometry (IO) using only Inertial Measurement Units (IMUs) offers a lightweight and cost-effective solution for Unmanned Aerial Vehicle (UAV) applications, yet existing learning-based IO models often fail to generalize to UAVs due to the highly dynamic and non-linear-flight patterns that differ from pedestrian motion. In this work, we identify that the conventional practice of transforming raw IMU data to global coordinates undermines the observability of critical kinematic information in UAVs. By preserving the body-frame representation, our method achieves substantial performance improvements, with a 66.7% average increase in accuracy across three datasets. Furthermore, explicitly encoding attitude information into the motion network results in an additional 23.8% improvement over prior results. Combined with a data-driven IMU correction model (AirIMU) and an uncertainty-aware Extended Kalman Filter (EKF), our approach ensures robust state estimation under aggressive UAV maneuvers without relying on external sensors or control inputs. Notably, our method also demonstrates strong generalizability to unseen data not included in the training set, underscoring its potential for real-world UAV applications.
@@
Learning-based Inertial Odometry
@@ I Introduction
Inertial Measurements Units (IMUs) are inexpensive and ubiquitous sensors that provide linear accelerations and angular velocities. Due to the size, weight, power, and cost (SWAP-C) constraints, IO based only on light and low-cost sensors is ideal for UAVs applications ranging from 3D mapping [1], exploration [2] to physical interaction [3]. Compared to exteroceptive sensors like vision and LiDAR, IMUs are unaffected by visual degradation factors such as motion blur [4] or dynamic object interruption [5], which are common in agile UAV flights [6]. Despite these advantages, current IO solutions struggle to accurately learn and adapt to the complex motion models of UAVs due to the inherent noise and bias of the IMU measurement, leading to large drift over time, particularly during agile maneuvers.
Figure 1: The upper figure shows quadrotor tilt dynamics: take-off, hovering, and landing maintain initial orientation; acceleration increases tilt for thrust; cruising holds steady tilt for constant velocity; and deceleration tilts oppositely to create drag. The lower figure depicts a simulated straight-line flight with constant yaw from the Pegasus Simulator [7]. The red dashed line tracks pitch angle adjustments, while black solid line shows velocity changes over time.
@@
Why is learning IO not directly applicable to drones?
@@
    We conduct the analysis of the extracted features from different representations using t-SNE and principal component analysis (PCA). The results show that body-frame representation is more expressive and observable. Simply changing the input representation can significantly improve IO accuracy by an average of 66.7%.
@@
    Grounded on this effective representation, we develop a learning-based IO framework AirIO that explicitly encodes the attitude information and fuses it with the body-frame IMU data to predict the velocity. Explicitly encoding the attitude information further improves accuracy by an additional 23.8%.
@@
    We further integrate an uncertainty-aware IMU preintegration model and the learned motion network into EKF for odometry estimation. Extensive real-world experiments validate our system, which outperforms existing IO algorithms without the need for additional sensors or control information. Our model also demonstrates generalizability to the unseen datasets that are not included in the training sets.
II Related Works
@@ II-A Model-based Inertial Odometry
Traditional model-based inertial odometry (IO) methods rely on kinematic motion models [14] to estimate relative motion and are often integrated with exteroceptive sensors such as cameras [16] and LiDAR [1]. While these algorithms achieve high accuracy, their dependence on external sensors makes them vulnerable to disturbances caused by agile motion or dynamic environments. To address the SWAP challenge of lightweight multirotor UAV navigation, Svacha et al. [17, 18] introduced a method that estimates tilt and velocity by fusing tachometer and IMU data. Although effective, this approach still rely on auxiliary sensor, limiting their applicability in environments where such sensors are unavailable or unreliable.
@@ II-B Learning Inertial Odometry
Recent advances in deep learning have sparked research in data-driven methods for learning velocity and displacement from inertial data [11, 19, 9, 20]. Approaches such as IONet [19], A2DIO [21], and NILoc [20] formulate inertial navigation as a sequential problem task by predicting relative position displacements using IMU data. Subsequent work has incorporated these learned displacements into EEKF [8, 22, 23] or batched optimization frameworks [24]. To ensure consistency in IMU measurements, many of these methods transform the input IMU data into global coordinates. For pedestrian navigation, RoNIN [9] proposed a heading-agnostic coordinate frame (HACF), aligning the accelerometer’s gravity vector with the z-axis and constraining rotation to the horizontal plane. This framework has been widely adopted for learning-based IO [8, 22] due to its effectiveness in simplifying pedestrian motion. Building on HACF, RIO [25] introduced rotation-equivalence data augmentation to self-supervise pose estimation after orientation alignment. Despite the success in pedestrian navigation and wheeled robots [9, 26], these representations are less effective for multirotor UAV motion, which require more sophisticated dynamic modeling. In this paper, we show that the commonly used global coordinate representations hinder neural networks from effectively capturing the dynamic complexities of UAVs.
@@ II-C Inertial Odometry for Multirotor UAVs
Existing learning-based IO methods have primarily focused on pedestrian datasets, limiting their generalizability to platforms with more demanding motion dynamics, such as UAVs. Approaches like AI-driven pre-processing and down-sampling of high-speed inertial data have been proposed for better state estimation [27]. DIDO [15] estimates the thrust of UAVs using tachometer data and addresses unmodeled forces by incorporating a neural network. Building on this, the IMO [6] incorporates thrust information and IMU data within an EKF framework, which implicitly captures the drone dynamics. However, IMO relies on additional control signals to capture the drone’s dynamic, as shown in Fig. 8, which may suffer from overfitting on training datasets. DIVE [28], a more recent work for UAV IO, encodes orientation alongside gravity-removed global-frame acceleration. While this representation improves the robustness of learning-based IO, our findings in Section. V indicate that it remains suboptimal for capturing UAV dynamics comprehensively.
@@ III Methodology
The goal of learning IO is to estimate the relative position transform (velocity) $\mathbf{v}$ from the IMU data $\{\mathbf{a}_{i},\mathbf{w}_{i}\}$ which is defined as $\mathbf{v}=\Psi(\{\mathbf{a}_{i},\mathbf{w}_{i}\})^{n}_{i=1}\in\mathbb{R}^{3}$ from the local body IMU frame. Most of the existing methods employ transforming the IMU frame to global frame or removing the gravity from the raw IMU data[9, 25, 28].
However, our study reveals that these widely used representations-such as global coordinate frames, HACF, and gravity-removed frames—limit the effectiveness of IMU feature extraction. Grounded on our analysis, our work consists of three parts: feature representation analysis, attitude encoding, and a tightly coupled EKF system.
We first investigate the feature representation of the learning IO in Section. III-A and Section. III-B, finding that the body coordinate frame is more representative than the other representation. Grounded on these findings, we propose AirIO which explicitly encode the drone’s pose to predict the body-frame velocity $\hat{\mathbf{v}}_{i}^{\mathcal{B}}$ and the corresponding uncertainty $\hat{\boldsymbol{\Sigma}}^{v}_{i}$ in Section. III-C. Lastly, in Section. III-D we build a tightly coupled EKF system that fuses the AirIO motion network with the learning-based IMU preintegration network, AirIMU [29], which jointly estimates the uncertainty of the IMU preintegration. Both modules predict their corresponding uncertainties, contributing to improved overall performance and robustness.
@@ III-A IMU Coordinate Frame
Canonical Body Coordinate Frame: The IMU accelerometer measures the net force per unit mass acting on the sensor. Due to the measurement mechanism, it measures not only the object actual accelerations due to motion $\mathbf{\dot{v}}^{\mathcal{B}}=\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}$ in the sensor’s local coordinate frame (usually aligned with object body frame), but also the constant acceleration due to Earth’s gravity $\mathbf{g}^{\mathcal{G}}$ directed toward the Earth’s center. Thus, the accelerometer measurement at timestamp $i$ as:
$\mathbf{a}_{i}^{\mathcal{B}}=\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}+\mathbf{R}_{i}^{\top}\mathbf{g}^{\mathcal{G}},$ (1)
where $m$ is the object mass, and $\mathbf{R}_{i}\in\mathbb{SO}(3)$ is the object attitude, which defines the rotation for vector from body frame $\mathcal{B}$ to the global frame $\mathcal{G}$. $(\cdot)^{\mathcal{F}}$ denotes quantity represented in frame $\mathcal{F}$.
Figure 2: (a) Existing learning IO transforms the IMU input to the global coordinate using ground-truth orientation. (b) We found that preserving the IMU data in the body coordinate, along with gravitational information, can significantly improve accuracy. (c) Grounded on our previous analysis, our AirIO model leverages the body-frame IMU and explicitly encodes the drone’s orientation to estimate the body-frame velocity of the drone.
Global Coordinate Frame: Many existing methods transform the IMU measurement from body frame $\mathcal{B}$ to HACF or global coordinate frame $\mathcal{G}$ by applying an estimated transform rotation $\hat{\mathbf{R}}_{i}$, e.g.
$\mathbf{a}_{i}^{\mathcal{G}}=\hat{\mathbf{R}}_{i}\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}+\hat{\mathbf{R}}_{i}\mathbf{R}_{i}^{\top}\mathbf{g}^{\mathcal{G}}$ (2)
@@
Notice that both (1) and (2) represents the same physical process. The key difference lies in how the attitude information is coupled with different components: in (1), it is coupled with the gravitational force, whereas in (2), it is coupled with the motion force. In (1), since the gravity acceleration $\mathbf{g}^{\mathcal{G}}$ is a constant, the the motion force and attitude form a linear combination. In contrast. in (2), the attitude couples with motion force in a nonlinear manner, with gravity acting as an additional constant that does not contribute extra useful information. Thus, the motion force and attitude are intuitively easier to estimate in (1), while (2) requires a more complex network structure for estimation.
@@ III-B Representation Analysis
Principal Component Analysis PCA is a statistical method that reduces the dimensionality of a large data set by replacing correlated variables with a smaller set of uncorrelated principal components. To analyze the expressivity of the input representation, we performed PCA on the feature latent space for different input coordinates. We concatenate the latent features from all samples and form a feature matrix. Then, we perform PCA on the latent feature matrix to analyze its underlying structure by computing the singular value of the normalized feature matrix in a numerically stable way. Each singular value represents the corresponding magnitude scaling of one principal feature component. We then quantify the cumulative explained variance (or the energy coverage) for the top $k$ principal components in PCA. Specifically, we evaluate on two datasets. One is the widely used UAV dataset Blackbird[30] collected by a custom UAV platform Blackbird for agile autonomous flight, which covers large UAV maneuvers. The other is our custom simulation dataset Pegasus, which we collected using the NVIDIA IssicSim simulator with Pegasus autopilot [7] also with diverse maneuvers but in the ideal environment without external disturbance, allowing us to work with cleaner data.
@@
Fig. 3 shows the cumulative energy coverage of the latent feature derived from two different IMU input representations: global and body frames. The red and black curves represent the energy coverage of the top $k$ principal components. Specially, for the Blackbird dataset, it shows that the top $k$ = 5 principal components in body-frame representation already cover 95% of the total energy. In contrast, global-frame representation requires the top $k$ = 8 principal components to achieve the same level of energy. Similarly, in the Pegasus datasets, the body-frame representation requires top $k$ = 15 principal components and the global frame require top $k$ = 40 principal components to cover 95% of the total energy. It clearly shows that models trained on the body representation input can achieve comparable performance with fewer features, highlighting the efficiency and expressivity of this representation in capturing the essential features. It also implies that models trained under body frame have better compressibility and can be designed more lightweight. Figure. 4 shows the comparison of the absolute trajectory error (ATE) for the models at different scales under both body and global frames. As the model size decreases, the prediction performance for model under body-frame representation degrades more smoothly and consistently yields lower errors, further showing the better compressibility and expressivity of the representation. More studies about the model compressibility can be found in Appendix. A-F.
Figure 4: The ATE for body-frame and global-frame model at different model sizes on Pegasus dataset (Left) and Blackbird dataset (Right). The text denotes the relative percentage increase in ATE.
t-SNE [31] We also perform t-SNE on Pegasus dataset, a dimensionality reduction approach that projects high-dimension data into a low-dimensional map to analyze the feature representation qualitatively. As shown in Fig. 5, based on the network prediction (here is the velocity), we color code each data point by the velocity magnitude from low to high, represented by blue to red. In the global coordinate frame, the points exhibit a highly entangled distribution and different velocity levels overlap, indicating poor separability and less representative encoding. In contrast, the feature distribution of the body coordinate frame is more structured and well-separated, where different velocity levels form more coherent clusters. This suggests that the body coordinate frame facilitates a more discriminative and effective representation of the latent features.
@@
In summary, the analysis from Sec.III-A and Sec. III-B clearly indicates that i). A good prior information will help for estimation as it couples with the motion force in linear or nonlinear manner as shown in (1)(2). ii). The nonlinear coupling between the motion force and attitude in global-frame representation complicates the estimation problem. The body-frame representation proves to be more expressive and discriminative, leading to light network for prediction.

### v2

@@ Abstract
Inertial odometry (IO) using only Inertial Measurement Units (IMUs) offers a lightweight and cost-effective solution for Unmanned Aerial Vehicle (UAV) applications, yet existing learning-based IO models often fail to generalize to UAVs due to the highly dynamic and non-linear-flight patterns that differ from pedestrian motion. In this work, we identify that the conventional practice of transforming raw IMU data to global coordinates undermines the observability of critical information in UAVs. By preserving the body-frame representation, our method achieves substantial performance improvements, with a 66.7% average increase in accuracy across three datasets. Furthermore, explicitly encoding attitude information into the motion network results in an additional 23.8% improvement over prior results. Combined with a data-driven IMU correction model (AirIMU) and an uncertainty-aware Extended Kalman Filter (EKF), our approach ensures robust state estimation under aggressive UAV maneuvers without relying on external sensors or control inputs. Notably, our method also demonstrates strong generalizability to unseen data, underscoring its potential for real-world UAV applications.
@@
Aerial Systems: Perception and Autonomy; Deep Learning Methods; Localization
@@ I Introduction
Inertial Measurements Units (IMUs) are inexpensive and ubiquitous sensors that provide linear accelerations and angular velocities. Due to the size, weight, power, and cost (SWAP-C) constraints, IO based only on light and low-cost sensors is ideal for Unmanned Aerial Vehicles (UAVs) applications ranging from 3D mapping [1], exploration [2] to physical interaction [3]. Compared to exteroceptive sensors like vision and LiDAR, IMUs are unaffected by visual degradation factors such as motion blur [4] or dynamic object interruption [5], which are common in agile UAV flights [6]. Despite these advantages, current IO solutions struggle to accurately adapt to the complex motion models of UAVs due to the IMU inherent noise and bias, leading to large drift over time, particularly during agile maneuvers.
@@
Why is learning IO not directly applicable to UAVs?
@@
    We conduct the analysis of the extracted features from different representations using t-Distributed Stochastic Neighbor Embedding (t-SNE) and principal component analysis (PCA). The results show that body-frame representation is more expressive and observable. Simply changing the input representation can significantly improve IO accuracy by an average of 66.7%.
@@
    Building on this, we develop AirIO, a learning-based IO that encodes the attitude information and fuses it with the body-frame IMU data for velocity prediction. This explicit encoding further improves accuracy by 23.8%. Additionally, we integrate an uncertainty-aware IMU preintegration model and a learned motion network into EKF for odometry estimation.
@@
    We validate our approach on extensive experiments. It outperforms existing IO algorithms without the need for additional sensors or control information. It also demonstrates generalizability to the unseen datasets.
II Related Work
@@ II-A Model-based Inertial Odometry
Traditional model-based inertial odometry (IO) methods rely on kinematic models [13] to estimate relative motion and are often integrated with exteroceptive sensors such as cameras [15] and LiDAR [1]. While these methods achieve high accuracy, their dependence on external sensors makes them vulnerable to disturbances caused by agile motion or dynamic environments. To address the SWAP challenge of lightweight UAV navigation, Ref [16] introduced a method that estimates tilt and velocity by fusing tachometer and IMU data. Although effective, this approach still relies on auxiliary sensors, limiting its applicability in environments where such sensors are unavailable or unreliable.
@@ II-B Learning Inertial Odometry
Recent advances in deep learning have sparked research in data-driven methods for learning velocity and displacement from inertial data [10, 17, 8, 18]. Approaches such as IONet [17], A2DIO [19], and NILoc [18] formulate inertial navigation as a sequential task by predicting relative position displacements using IMU data. Subsequent work has incorporated these learned displacements into EKF [7, 20, 21] or batched optimization frameworks [22]. To ensure consistency in IMU measurements, many of these methods transform the input IMU data into global coordinates using ground-truth orientation. For pedestrian navigation, RoNIN [8] proposed a Heading-Agnostic Coordinate Frame (HACF), aligning the accelerometer’s gravity vector with the z-axis and constraining rotation to the horizontal plane. This framework has been widely adopted for learning-based IO [7, 20] due to its effectiveness in simplifying pedestrian motion. Building on HACF, RIO [23] introduced rotation-equivalence data augmentation to self-supervise pose estimation after orientation alignment. Since pedestrians and wheeled vehicles predominantly move along the horizontal plane, HACF effectively removes yaw dependency from the representation, thereby reducing learning complexity and improving generalizability. Despite the success in pedestrian navigation and wheeled robots [8, 24], these representations are less effective for multirotor UAV motion, which requires more sophisticated dynamic modeling. In this paper, we show that the commonly used global coordinate representations hinder neural networks from effectively capturing the dynamic complexities of UAVs.
@@ II-C Inertial Odometry for Multirotor UAVs
Existing learning-based IO methods have primarily focused on pedestrian datasets, limiting their generalizability to platforms with more demanding motion dynamics, such as UAVs. Approaches like AI-driven pre-processing and down-sampling of high-speed inertial data have been proposed for better state estimation [25]. Due to its simplicity, most of the existing methods continue adopting the global coordinate frame, such as [14, 6, 26]. DIDO [14] estimates the thrust of UAVs using tachometer data and addresses unmodeled forces by incorporating a neural network. Building on this, the IMO [6] incorporates thrust information and IMU data within an EKF framework, which implicitly captures the drone dynamics. However, IMO relies on additional control signals to capture the drone’s dynamics, as shown in Fig. 7, which may suffer from overfitting on training datasets. DIVE [26], a more recent work for UAV IO, encodes orientation alongside gravity-removed global-frame acceleration. While this representation improves the robustness of learning-based IO, our findings in Section. V indicate that it remains suboptimal for capturing UAV dynamics comprehensively.
@@ III Methodology
The goal of learning IO is to estimate relative position transform (velocity) from the IMU data in the local body IMU frame. Most existing methods employ transforming the IMU frame to global frame or removing gravity from the raw IMU data[8, 23, 26]. Our study reveals that these widely used representations, such as global coordinate frames, HACF, and gravity-removed frames, limit the effectiveness of IMU feature extraction. Thus, this section proposes a different feature representation and the corresponding state estimation design. It consists of three parts: feature representation analysis, attitude-encoded network design, and a tightly coupled EKF system.
@@ III-A IMU Coordinate Frame
Canonical Body Coordinate Frame: The IMU accelerometer measures the net force per unit mass acting on the sensor. Due to the measurement mechanism, it measures not only the object actual accelerations due to motion ${}^{\mathcal{B}}\mathbf{\dot{v}}=\frac{{}^{\mathcal{B}}\mathbf{F}_{{net}_{i}}}{m}$ in the sensor’s local coordinate frame (usually aligned with object body frame), but also the constant acceleration due to Earth’s gravity ${}^{\mathcal{G}}\mathbf{g}$ directed toward the Earth’s center. Thus, the accelerometer measurement at timestamp $i$ as:
${}^{\mathcal{B}}\mathbf{a}_{i}=\frac{{}^{\mathcal{B}}\mathbf{F}_{{net}_{i}}}{m}+\mathbf{R}_{i}^{\top}{}^{\mathcal{G}}\mathbf{g},$ (1)
where $m$ is the object mass, and $\mathbf{R}_{i}\in\mathbb{SO}(3)$ is the rotation for vector from body frame $\mathcal{B}$ to the global frame $\mathcal{G}$. ${}^{\mathcal{F}}(\cdot)$ denotes quantity represented in frame $\mathcal{F}$.
Figure 1: (a) Existing learning IO usually transforms the IMU input to the global coordinate using ground-truth orientation. (b) We found that preserving the IMU data in the body coordinate, along with gravitational information, can significantly improve accuracy. (c) Our AirIO model leverages the body-frame IMU and explicitly encodes the UAV’s orientation to estimate the body-frame velocity.
Global Coordinate Frame: Many existing methods transform the IMU data from body frame $\mathcal{B}$ to HACF or global frame $\mathcal{G}$ by applying an estimated transform rotation $\hat{\mathbf{R}}_{i}$, e.g.
${}^{\mathcal{G}}\mathbf{a}_{i}=\hat{\mathbf{R}}_{i}\frac{{}^{\mathcal{B}}\mathbf{F}_{{net}_{i}}}{m}+\hat{\mathbf{R}}_{i}\mathbf{R}_{i}^{\top}{}^{\mathcal{G}}\mathbf{g}$ (2)
@@
Notice that both (1) and (2) represents the same physical process. The key difference lies in how the attitude information is coupled with different components: in (1), it is coupled with the gravitational force, whereas in (2), it is coupled with the motion force. In (1), since the gravity acceleration ${}^{\mathcal{G}}\mathbf{g}$ is a constant, the the motion force and attitude form a linear combination. In contrast. in (2), the attitude couples with motion force in a nonlinear manner, with gravity acting as an additional constant that does not contribute extra useful information. Thus, the motion force and attitude are intuitively easier to estimate in (1), while (2) requires a more complex network structure for estimation.
@@ III-B Representation Analysis
Principal Component Analysis PCA is a statistical method that reduces the dimensionality of a data set by replacing correlated variables with a smaller set of uncorrelated principal components. To analyze the expressivity of the input representation, we performed PCA on the feature latent space for different input coordinates. We concatenate the latent features from all samples to form a feature matrix and perform PCA on the latent feature matrix to analyze its underlying structure by computing the singular value of the normalized feature matrix in a numerically stable way. Each singular value represents the corresponding magnitude scaling of one principal feature component. We then quantify the cumulative explained variance (or the energy coverage) for the top $k$ principal components in PCA. Specifically, we evaluate on two datasets: the widely used UAV dataset Blackbird[27] collected for agile autonomous flight, which covers large UAV maneuvers, and our custom simulation dataset Pegasus, which which was collected in the NVIDIA IssacSim simulator with Pegasus autopilot [28], featuring diverse maneuvers in an ideal environment without external disturbance, allowing us to work with cleaner data.
@@
Fig. 2 shows the cumulative energy coverage of the latent feature derived from two different IMU input representations: global and body frames. The red and black curves represent the energy coverage of the top $k$ principal components. Specifically, for the Blackbird dataset, it shows that the top $k$ = 5 principal components in body-frame representation already cover 95% of the total energy. In contrast, global-frame representation requires the top $k$ = 8 principal components to achieve the same level of energy. Similarly, in the Pegasus datasets, the body-frame representation requires top $k$ = 15 principal components and the global frame require top $k$ = 40 principal components to cover 95% of the total energy. It clearly shows that models trained on the body representation input can achieve comparable performance with fewer features, highlighting the efficiency and expressivity of this representation in capturing the essential features. It also implies that models trained under body frame have better compressibility and can be designed more lightweight. Fig. 3 shows the comparison of the absolute translation error (ATE) for the models at different scales under both body and global frames. As the model size decreases, the prediction performance for model under body-frame representation degrades more smoothly and consistently yields lower errors. See Appendix. A-G for more studies on the model compressibility.
Figure 3: The ATE for body-frame and global-frame models at different model sizes on Pegasus dataset (Left) and Blackbird dataset (Right). The text denotes the relative percentage increase in ATE.
t-SNE [29] We also perform t-SNE on Pegasus dataset, a dimensionality reduction approach that projects high-dimension data into a low-dimensional map to analyze the feature representation qualitatively. As shown in Fig. 4, based on the network prediction (here is the velocity), we color code each data point by the velocity magnitude from low to high. In the global frame, the points exhibit a highly entangled distribution and different velocity levels overlap, indicating poor separability and less representative encoding. In contrast, the feature distribution of the body coordinate frame is well-separated, where different velocity levels form more coherent clusters. This suggests that the body frame facilitates a more discriminative and effective representation of latent features.
@@
In summary, the analysis from Section. III-A and Section. III-B indicates that i). Good prior information will help for the estimation since it couples with the motion force in a linear or nonlinear manner, as shown in (1)(2). ii). The nonlinear coupling between the motion force and attitude in the global-frame representation complicates the estimation problem. Body-frame representation proves to be more expressive and discriminative, leading to a light network for prediction. Notice that while the analysis is indeed applicable to general robots via standard IMU equations, UAVs present a compelling test case due to their agile motion. Unlike many mobile or legged robots with mainly yaw-dominant movement, UAVs often exhibit significant roll and pitch maneuvers, resulting in more pronounced impact of frame representation difference.